OpenStudy (anonymous):

solve the following; 1/6x-3/10x=1/6

OpenStudy (anonymous):

(1/6)x-(3/10)x=1/6 ,mult both sides by 6 we get x-1.8x=1 -(4/5)x=1 x=-5/4